OpenStudy (anonymous):

cot^2x-(1/sin^2x) find the simplest expression

OpenStudy (anonymous):

\[ \cot^2(x) = \frac{\cos^2(x)}{\sin^2(x)} \]

OpenStudy (anonymous):

\[\frac{ \cos^2x }{ \sin^2x }-\frac{ 1 }{ \sin^2x }=\frac{ \cos^2x -1}{ \sin^2x }=-\frac{ (1-\cos^2x) }{ \sin^2x }=-\frac{ \sin^2x }{ \sin^2x }=-1\] @rl88

OpenStudy (anonymous):

\[ x\ne \pi n,\quad n\in \mathbb{Z} \]

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

Medal? lol @rl88

OpenStudy (anonymous):

umm...suree...i think i gave you one?? wasnt too sure how to do it!